5. Find the general solution to the following non-homogeneous differential equation. x – 2x + x = t2 +t+1
Answer #1

al-2x+x=4²+t+1 To find C.F. (02-20+1) =k+t+1, 4.DE Let Ho be a trial solution, then auxilliary equation is m²- 2m+1=0 => (m-A - 2.372.1 21 => A6+2 =) 2) A-4 21 2) A = 5 xp =.5d3t+t2 :: General solutionis X=(@+ext)e+ 5+3t+ť? where a, la are arbitrar